# Shortcuts Everywhere and Nowhere: Exploring Multi-Trigger Backdoor Attacks

### v2

@@ Abstract
Backdoor attacks have become a significant threat to the pre-training and deployment of deep neural networks (DNNs). Although numerous methods for detecting and mitigating backdoor attacks have been proposed, most rely on identifying and eliminating the “shortcut” created by the backdoor, which links a specific source class to a target class. However, these approaches can be easily circumvented by designing multiple backdoor triggers that create shortcuts everywhere and therefore nowhere specific. In this study, we explore the concept of Multi-Trigger Backdoor Attacks (MTBAs), where multiple adversaries leverage different types of triggers to poison the same dataset. By proposing and investigating three types of multi-trigger attacks including parallel, sequential, and hybrid attacks, we demonstrate that 1) multiple triggers can coexist, overwrite, or cross-activate one another, and 2) MTBAs easily break the prevalent shortcut assumption underlying most existing backdoor detection/removal methods, rendering them ineffective. Given the security risk posed by MTBAs, we have created a multi-trigger backdoor poisoning dataset to facilitate future research on detecting and mitigating these attacks, and we also discuss potential defense strategies against MTBAs.
Machine Learning, ICML
1 Introduction
Deep neural networks (DNNs) have become the standard models for tasks in computer vision (He et al., 2016; Dosovitskiy et al., 2021), natural language processing (Devlin et al., 2019; Mann et al., 2020), and speech recognition (Chan et al., 2016). However, research has demonstrated that DNNs are susceptible to backdoor attacks (Gu et al., 2017; Chen et al., 2017; Li et al., 2020), where stealthy triggers are embedded into the target models during training by poisoning a small subset of the training data or altering the training process. In image classification, this typically involves adding a fixed, carefully crafted trigger pattern to a few training images that the adversary can access. The objective of a backdoor attack is to manipulate the model into producing a specific output chosen by the adversary whenever the trigger pattern appears in a test input. With the increasing use of large models pre-trained on unsupervised web data or those provided by untrusted sources, concerns about the backdoor vulnerability of these models have grown, particularly when they are used in safety-critical applications.
@@ 1 Introduction
Given the significant security threat posed by backdoor attacks on neural networks, a variety of backdoor detection and removal techniques have been developed. Some prominent methods include (Wang et al., 2019; Li et al., 2021c; Wu & Wang, 2021; Li et al., 2023). These methods generally operate under the assumption that backdoor triggers are unknown in practice. As a result, they focus on identifying backdoors by detecting the presence of “shortcuts” between a specific source class and a target class, and then mitigating the threat by eliminating these shortcuts. However, in this work, we systematically demonstrate that such approaches can be effectively bypassed by implementing multiple backdoor triggers that create shortcuts in all directions, rendering them indistinguishable and making it challenging to target any specific one for removal. This strategy exposes a critical vulnerability in existing backdoor defenses, highlighting the need for more robust and comprehensive protection mechanisms.
In this work, we extend traditional single-trigger backdoor attacks to multi-trigger backdoor attacks. Note that we do not call them multi-adversary attacks as our focus is on the triggers, demonstrating that even a single adversary can simultaneously employ multiple distinct triggers to poison a dataset. We explore three distinct implementation strategies for multi-trigger backdoor attacks: parallel, sequential, and hybrid, where the latter blends multiple trigger patterns into a single, more potent super trigger pattern. Furthermore, we conduct a thorough analysis of 10 different existing triggers, uncovering their coexistence, overwriting, and cross-activation effects. Our findings reveal that: 1) different backdoor triggers can coexist within the same dataset, whether injected in parallel or sequentially; 2) certain triggers can overwrite others even at minimal poisoning rates (as low as 0.2% or 0.5%); 3) one trigger can cross-activate another in sequential attack settings; and 4) different trigger patterns can be combined into a hybrid trigger to enhance attack efficacy. Figure 1 summarizes some of the experimental results. These insights highlight the heightened security risks posed by the presence of multiple backdoor adversaries.
@@
Next, we systematically evaluate existing backdoor defense methods since they have not been adequately evaluated against multi-trigger attacks. This gap in evaluation can lead to overly optimistic perceptions of backdoor robustness and a false sense of security. Our findings confirm this vulnerability: multi-trigger backdoor attacks can severely undermine the effectiveness of current state-of-the-art backdoor detection and removal techniques. Specifically, 1) widely used detection methods like Neural Cleanse (NC) (Wang et al., 2019) struggle to identify multi-trigger attacks, and 2) mainstream backdoor removal techniques such as fine-tuning (Liu et al., 2018a), Neural Attention Distillation (NAD) (Li et al., 2021c), Adversarial Neuron Pruning (ANP) (Wu & Wang, 2021), Reconstructive Neuron Pruning (RNP) (Li et al., 2023), and Anti-Backdoor Learning (ABL) fail to effectively purify models compromised by multiple triggers. The challenge of identifying multiple triggers (or the correct sequence to defend against them) is compounded by the difficulty of determining the exact number of triggers present, making defense against multi-trigger attacks extremely difficult.
Our main contributions are as follows.
@@
2 Related Work
Backdoor Attack
A backdoor attack designs a trigger pattern to achieve a high attack success rate while being stealthy to humans. Existing backdoor triggers can be categorized into dataset-wise vs. sample-wise triggers. A dataset-wise trigger applies the same trigger pattern to all poisoned samples while a sample-wise trigger poisons each sample with a unique trigger pattern. Examples of dataset-wise triggers include one pixel (Tran et al., 2018), a checkerboard pattern (Gu et al., 2017), a global pattern like Gaussian noise (Chen et al., 2017), background reflection (Liu et al., 2020), to name a few. Compared to dataset-wise triggers, sample-wise triggers can be more complex (and stealthy) as they are often optimized by additional techniques. For instance, generative model-based backdoor attacks (Lin et al., 2020; Nguyen & Tran, 2020; Cheng et al., 2021; Li et al., 2021a) generate backdoor trigger patterns adaptively based on different input samples. Regardless of the diverse trigger patterns, all existing backdoor attacks follow a single-trigger attack setting, where there exists only one adversary and one single type of trigger (Li et al., 2022). Such a restricted setting limits the exploration of backdoor attacks in more realistic settings where multiple adversaries could attack the same dataset. It also raises the question as to whether these attacks can coexist and be simultaneously effective when leveraged by different adversaries to poison the same dataset. In this paper, we explore such an aspect of a diverse set of existing triggers under a more realistic setting: multi-trigger attack setting where there are multiple adversaries and types of triggers.
Backdoor Defense
@@ 2 Related Work
Existing backdoor defenses can be categorized into backdoor detection and backdoor removal (or mitigation) methods. Detection methods identify whether a given model has been backdoored by a backdoor attack (Chen et al., 2019; Wang et al., 2019; Guo et al., 2019; Xu et al., 2021) or whether a sample contains a backdoor trigger (Tran et al., 2018; Chen et al., 2019; Gao et al., 2019; Zeng et al., 2021; Huang et al., 2023).
@@
Removal methods aim to eliminate the backdoor trigger (if it exists) from the model without affecting its functionality. This can be done during the training process via anti-backdoor learning (Li et al., 2021b; Huang et al., 2022), or later on via fine-tuning, fine-pruning (Liu et al., 2018a), distillation (Li et al., 2021c), adversarial pruning (Wu & Wang, 2021), channel Lipschitzness based pruning (Zheng et al., 2022), or reconstructive pruning (Li et al., 2023). However, all of these defenses were developed for single-trigger attacks, thus facing high uncertainty when applied to defend multi-trigger attacks. In this paper, we run extensive experiments to reexamine the effectiveness of these defenses to multi-trigger backdoor attacks.
3 Multi-Trigger Backdoor Attacks
@@
3.1 Threat Model and Problem Definition
Threat Model
Our threat model introduces 3 poisoning strategies for multi-trigger attacks: parallel, sequential, and hybrid, covering scenarios involving both independent and collusive adversaries. Specifically, there exist one or more adversaries where each adversary can access a few training samples but none of them can manipulate the training procedure. The adversaries can be independent or collusive, with independent adversaries poisoning different training samples with different types of triggers, while collusive adversaries could poison the same training samples in sequential order by stacking one trigger after another. There could also exist a super adversary that launches the above attacks altogether or combines multiple triggers into a hybrid trigger.
Definition of Backdoor Attack
We focus on image classification tasks. Given a clean training dataset ${\mathcal{D}}=\{({\bm{x}}_{i},y_{i})\}_{i=1}^{N}$, where ${\bm{x}}_{i}\in{\mathcal{X}}$ represents a training image and $y_{i}\in{\mathcal{Y}}$ is its label. A backdoor adversary generates backdoor examples with a triggering function $tr:{\mathcal{X}}\rightarrow{\mathcal{X}}$. For each clean sample, it maps the sample ${\bm{x}}$ into a backdoor sample ${\bm{x}}_{b}$, i.e., ${\bm{x}}_{b}=\operatorname{tr}({\bm{x}})$ and modifies its label to a backdoor target label $y_{b}$. To ensure stealthiness and efficacy for backdoor injection, the adversary randomly chooses a few training samples to poison, which creates a set of backdoor samples ${\mathcal{D}}_{b}=\{({\bm{x}}_{b},y_{b})\}$. As we mentioned earlier, most existing attacks only inject a single type of backdoor trigger into the training data, meaning there is only one triggering function $\operatorname{tr}(\cdot)$. We call these attacks Single-Trigger Backdoor Attacks (STBAs). The final poisoned training dataset can be denoted as $\hat{{\mathcal{D}}}={\mathcal{D}}_{c}\cup{\mathcal{D}}_{b}$, where ${\mathcal{D}}_{c}=\{({\bm{x}}_{c},y_{c})\}$ represents clean samples and clean labels, while ${\mathcal{D}}_{b}=\{({\bm{x}}_{b},y_{b})\}$ represents backdoor samples and backdoor labels. The poisoning rate is defined as $\alpha=|{\mathcal{D}}_{b}|/|\hat{{\mathcal{D}}}|$. Training a model on $\hat{{\mathcal{D}}}$ is solving the following optimization problem:
@@ 3.1 Threat Model and Problem Definition
Definition of Multi-Trigger Backdoor Attack
We call backdoor attacks that utilize multiple types of triggers (possibly from one or more adversaries) to attack the same dataset termed Multi-Trigger Backdoor Attacks (MTBAs). Figure LABEL:fig:overview illustrates the idea of MTBA. Note that we did not call these attacks multi-adversary attacks as even one adversary could implement different types of triggers. The MTBA problem can also be formulated as Eq. (1), but with a slightly more complex poisoning set that contains multiple triggers:
@@
where ${\bm{x}}^{k}_{b}=\operatorname{tr}_{k}({\bm{x}})$ is a poisoned sample by the $k$-th trigger $\operatorname{tr}_{k}(\cdot)$, and ${\mathcal{D}}^{k}_{b}$ is the subset of poisoned samples by trigger $\operatorname{tr}_{k}(\cdot)$ for overall $m$ triggers. For multi-trigger attacks, the poisoned dataset becomes $\hat{{\mathcal{D}}}={\mathcal{D}}_{c}\cup{\mathcal{D}}_{b}$ with a poisoning rate of $\alpha=|{\mathcal{D}}_{b}|/|\hat{{\mathcal{D}}}|=\sum_{k=1}^{m}|{\mathcal{D}}^{k}_{b}|/|\hat{{\mathcal{D}}}|$.
3.2 Poisoning Strategies
@@ 3.2.2 Sequential Poisoning
3.2.3 Hybrid-trigger Poisoning
@@
Candidate Triggers
Gathering a representative and meaningful set of triggers ${\mathcal{T}}$ is crucial for our study. By investigating the current literature, we meticulously selected 10 types of triggers used in mainstream backdoor attacks as our candidate triggers, which include both dataset-wise triggers (such as one pixel, a checkerboard pattern, and adversarial perturbation) and sample-wise triggers (such as input-aware pattern and image deformation). A detailed list of the studied triggers can be found in Section 4.1.
3.3 Label Modification
In the context of multi-trigger backdoor attacks, different adversaries may share the same backdoor target if it is of general interest, or they may have entirely different (or random) target labels. Following prior works (Gu et al., 2017; Xue et al., 2020), we explore the following three label modification strategies, including All2One, All2All, and All2Random (see Figure 3).

### v4

@@ Abstract
Backdoor attacks have become a significant threat to the pre-training and deployment of deep neural networks (DNNs). Although numerous methods for detecting and mitigating backdoor attacks have been proposed, most rely on identifying and eliminating the “shortcut” created by the backdoor, which links a specific source class to a target class. However, these approaches can be easily circumvented by designing multiple backdoor triggers that create shortcuts everywhere and therefore nowhere specific. In this study, we explore the concept of Multi-Trigger Backdoor Attacks (MTBAs), where multiple adversaries leverage different types of triggers to poison the same dataset. By proposing and investigating three types of multi-trigger attacks including parallel, sequential, and hybrid attacks, we demonstrate that 1) multiple triggers can coexist, overwrite, or cross-activate one another, and 2) MTBAs easily break the prevalent shortcut assumption underlying most existing backdoor detection/removal methods, rendering them ineffective. Given the security risk posed by MTBAs, we have created a multi-trigger backdoor poisoning dataset to facilitate future research on detecting and mitigating these attacks, and we also discuss potential defense strategies against MTBAs. Our code is available at https://github.com/bboylyg/Multi-Trigger-Backdoor-Attacks.
I Introduction
Deep neural networks (DNNs) have become the standard models for tasks in computer vision [1, 2], natural language processing [3, 4], and speech recognition [5]. However, research has demonstrated that DNNs are susceptible to backdoor attacks [6, 7, 8], where stealthy triggers are embedded into the target models during training by poisoning a small subset of the training data or altering the training process. In image classification, this typically involves adding a fixed, carefully crafted trigger pattern to a few training images that the adversary can access. The objective of a backdoor attack is to manipulate the model into producing a specific output chosen by the adversary whenever the trigger pattern appears in a test input. With the increasing use of large models pre-trained on unsupervised web data or those provided by untrusted sources, concerns about the backdoor vulnerability of these models have grown, particularly when they are used in safety-critical applications.
@@ I Introduction
Given the significant security threat posed by backdoor attacks on neural networks, a variety of backdoor detection and removal techniques have been developed. Some prominent methods include [9, 10, 11, 12]. These methods generally operate under the assumption that backdoor triggers are unknown in practice. As a result, they focus on identifying backdoors by detecting the presence of “shortcuts” between a specific source class and a target class, and then mitigating the threat by eliminating these shortcuts. However, in this work, we systematically demonstrate that such approaches can be effectively bypassed by implementing multiple backdoor triggers that create shortcuts in all directions, rendering them indistinguishable and making it challenging to target any specific one for removal. This strategy exposes a critical vulnerability in existing backdoor defenses, highlighting the need for more robust and comprehensive protection mechanisms.
In this work, we extend traditional single-trigger backdoor attacks to multi-trigger backdoor attacks. Note that we do not call them multi-adversary attacks as our focus is on the triggers, demonstrating that even a single adversary can simultaneously employ multiple distinct triggers to poison a dataset. We explore three distinct implementation strategies for multi-trigger backdoor attacks: parallel, sequential, and hybrid, where the latter blends multiple trigger patterns into a single, more potent super trigger pattern. Furthermore, we conduct a thorough analysis of 10 different existing triggers, uncovering their coexistence, overwriting, and cross-activation effects. Our findings reveal that: 1) different backdoor triggers can coexist within the same dataset, whether injected in parallel or sequentially; 2) certain triggers can overwrite others even at minimal poisoning rates (as low as 0.2% or 0.5%); 3) one trigger can cross-activate another in sequential attack settings; and 4) different trigger patterns can be combined into a hybrid trigger to enhance attack efficacy. Fig. 1 summarizes some of the experimental results. These insights highlight the heightened security risks posed by the presence of multiple backdoor adversaries.
@@
Next, we systematically evaluate existing backdoor defense methods since they have not been adequately evaluated against multi-trigger attacks. This gap in evaluation can lead to overly optimistic perceptions of backdoor robustness and a false sense of security. Our findings confirm this vulnerability: multi-trigger backdoor attacks can severely undermine the effectiveness of current state-of-the-art backdoor detection and removal techniques. Specifically, 1) widely used detection methods like Neural Cleanse (NC) [9] struggle to identify multi-trigger attacks, and 2) mainstream backdoor removal techniques such as fine-tuning [13], Neural Attention Distillation (NAD) [10], Adversarial Neuron Pruning (ANP) [11], Reconstructive Neuron Pruning (RNP) [12], and Anti-Backdoor Model (ABM) [7] fail to effectively purify models compromised by multiple triggers. The challenge of identifying multiple triggers (or the correct sequence to defend against them) is compounded by the difficulty of determining the exact number of triggers present, making defense against multi-trigger attacks extremely difficult.
Our main contributions are as follows:
@@
It is essential to emphasize that our work represents a significant attempt at designing MTBAs to simulate a more realistic scenario. The introduction of MTBAs itself undoubtedly provides a novel attack paradigm, offering new perspectives and design strategies for future research. We believe that MTBAs are an important leap compared to the existing single-trigger setting, and hope more researches could focus on potential threat of MTBAs.
The remainder of this paper is organized as follows: In section II, we briefly review the related work on backdoor attacks and defenses. Following that, we present the threat model and formally define the problem studied in this paper. In Section IV, we introduce the technical details of our proposed multi-trigger backdoor attack strategies. Section V presents extensive experiments on multiple benchmark datasets, validating the effectiveness and impact of our method. In Section VI, we re-evaluate the performance of existing backdoor defense techniques against multi-trigger backdoor attacks and analyze the challenges they face. Finally, Section VII concludes the paper by summarizing our findings and discussing future research directions. We hope our work provides new insights into backdoor threats and motivates further research on more robust defense mechanisms.
TABLE I: Comparison of existing multi-trigger backdoor attack studies and our MTBAs framework.
Paper & Year Trigger Type(s) Poisoning Mode Evaluation Scope Model Architecture Defense Evaluation Interaction Analysis Xue et al., TDSC 2022 [14] Static triggers (1–2) Dataset-level, single shot $\leq$2 triggers, 1 dataset CNN only (CIFAR-10) Basic (All2One) Partial overwrite Xue et al., Appl Intell 2024 [15] Imperceptible multi-channel Sample-wise, single shot $\leq$2 triggers, 1 dataset CNN only No None Singh et al., AISP 2023 [16] Physical triggers, same target Real-world setup, single shot Physical attacks, 1 at a time CNN only No None Hou et al., TCSVT 2024 [17] Multi-trigger, multi-target Single shot $\leq$2 triggers, 1 dataset CNN only Limited None Sun et al., ICSIP 2024 [18] Frequency-domain, invisible Multi-trigger, single sample $\leq$2 triggers, 1 dataset CNN only No None MTBA (Ours) 10 diverse (static + sample-wise) Parallel / Sequential / Hybrid 10 triggers $\times$ 2 datasets (CIFAR-10, ImageNet) CNN + ViT (4 models) Yes (8 defenses) Yes (Coexistence, Overwrite, Cross-activation)
II Related Work
II-A Backdoor Attacks
A backdoor attack implants a trigger pattern into a model so that the model predicts an adversary-chosen label when the trigger is present, while maintaining normal performance on clean data. Existing backdoor triggers can be broadly categorized into dataset-wise and sample-wise triggers. A dataset-wise trigger applies the same perturbation to all poisoned samples, such as a single pixel [19], a checkerboard pattern [6], global Gaussian noise [7], or background reflection [20]. In contrast, sample-wise triggers generate diverse perturbations for each sample, often through optimization or generative models, e.g., composite backdoors [21], input-aware triggers [22], deep generative triggers [23], or invisible triggers [24].
Regardless of their diversity, most prior works adopt a single-trigger attack setting, where there exists only one adversary and one type of trigger [25]. This restricted setting limits the understanding of how backdoors behave in realistic environments, where multiple adversaries may poison the same dataset with different triggers. This naturally raises the question of whether such attacks can coexist, compete, or interact when injected simultaneously.
Recent studies have explored multi-trigger attacks. For example, Xue et al.[14], Singh et al.[16], Hou et al.[17], and Sun et al.[18] investigate scenarios with more than one trigger. However, these works remain limited—typically considering only two or three handcrafted triggers, restricted to CNN architectures, or focusing on narrow evaluation settings (e.g., All-to-One attacks on images).
Comparison with our MTBAs. Table I contrasts existing multi-trigger backdoor attacks with our proposed MTBAs. The contribution of MTBAs extends beyond simply introducing multiple triggers: 1) Trigger types. Prior work typically considers only a few handcrafted triggers (e.g., 1–2 visual patches or frequency-domain perturbations). In contrast, our benchmark incorporates 10 diverse triggers, including both static and sample-wise variants; 2) Attack modes. Earlier studies largely rely on one-shot settings. Our framework introduces three poisoning strategies—parallel, sequential, and hybrid—that more realistically simulate coexistence or collusion between adversaries; 3) Evaluation scope. Most prior works tested only a few triggers on a single CNN model. We systematically evaluate 10 triggers $\times$ 4 models (2 CNNs and 2 ViTs) $\times$ 2 datasets (CIFAR-10 and an ImageNet subset), providing a much broader and more rigorous scope; 4) Model architectures. Previous studies restrict evaluations to CNNs, while our work includes both CNNs and Vision Transformers (ViTs), highlighting vulnerabilities across architectures; 5) Defense evaluation. Earlier multi-trigger studies rarely evaluated defenses systematically. We assess eight mainstream defenses (four detection and four removal) under All-to-One, All-to-All, and All-to-Random settings, revealing their limitations; 6) Interaction analysis. To our knowledge, no prior study investigates how triggers interact. Our work is the first to analyze coexistence, overwriting, and cross-activation effects, showing that novel multi-trigger attack threats for evaluating real-world risks.
II-B Backdoor Defense
@@
Existing backdoor defenses can be categorized into backdoor detection and backdoor removal (or mitigation) methods. Detection methods identify whether a given model has been backdoored by a backdoor attack [26, 9, 27, 28] or whether a sample contains a backdoor trigger [19, 26, 29, 30, 31].
@@
Removal methods aim to eliminate the backdoor trigger (if it exists) from the model without affecting its functionality. This can be done during the training process via anti-backdoor learning [32, 33], or later on via fine-tuning, fine-pruning [13], distillation [10], adversarial pruning [11], channel Lipschitzness based pruning [34], or reconstructive pruning [12]. The Anti-Backdoor Model (ABM) [35] is a representative example, exploiting trigger interaction properties to develop a non-invasive defense mechanism. However, all of these defenses were developed for single-trigger attacks, thus facing high uncertainty when applied to defend multi-trigger attacks. In this paper, we run extensive experiments to reexamine the effectiveness of these defenses to multi-trigger backdoor attacks.
III Multi-Trigger Backdoor Attacks
@@
III-A Threat Model
Our threat model introduces 3 poisoning strategies for multi-trigger attacks: parallel, sequential, and hybrid, covering scenarios involving both independent and collusive adversaries:
  • Parallel Attack: In this scenario, independent adversaries act simultaneously but target different subsets of the training data. Each adversary introduces a distinct trigger or a set of triggers into their respective data portion.
  • Sequential Attacks: In this strategy, adversaries poison the same subset of training samples in a sequential order. Each adversary introduces one or more triggers into these samples, stacking them one after another.
  • Hybrid Attacks: In the hybrid attack scenario, adversaries operate in a collusive manner, injecting multiple backdoor triggers into overlapping datasets, leading to a situation where the model is simultaneously influenced by a combination of parallel and sequential triggers.
Specifically, in parallel and sequential MTBA, we validate the simultaneous existence of multiple independent attackers (up to 10) and analyze the case of independent attacks (as detailed in Sections IV-A and IV-B). In hybrid MTBA, we consider scenarios where multiple attackers collaborate to attack the same data samples (as detailed in Section IV-C). This diversity allows us to gain a comprehensive understanding of how different types of triggers can coexist and interfere with each other, which can inspire backdoor research in the future.
III-B Definition of Backdoor Attack
[…]
@@ III-B Definition of Backdoor Attack
III-C Definition of Multi-Trigger Backdoor Attack
We call backdoor attacks that utilize multiple types of triggers (possibly from one or more adversaries) to attack the same dataset termed Multi-Trigger Backdoor Attacks (MTBAs). Fig. 2 illustrates the idea of MTBA. Note that we did not call these attacks multi-adversary attacks as even one adversary could implement different types of triggers. The MTBA problem can also be formulated as Eq. (1), but with a slightly more complex poisoning set that contains multiple triggers:
@@ III-C Definition of Multi-Trigger Backdoor Attack
[…]
IV Poisoning Strategies of MTBAs
@@ IV-B Sequential Poisoning
IV-C Hybrid-trigger Poisoning
@@
Candidate Triggers. Gathering a representative and meaningful set of triggers $\mathcal{T}$} is crucial for our study. By investigating the current literature, we meticulously selected 10 types of triggers used in mainstream backdoor attacks as our candidate triggers, which include both dataset-wise triggers (such as one pixel, a checkerboard pattern, and adversarial perturbation) and sample-wise triggers (such as input-aware pattern and image deformation). A detailed list of the studied triggers can be found in Section V-A.
IV-D Label Modification
In the context of multi-trigger backdoor attacks, different adversaries may share the same backdoor target if it is of general interest, or they may have entirely different (or random) target labels. Following prior works [6, 14], we explore the following three label modification strategies, including All2One, All2All, and All2Random (see Fig. 3).